# Mobility Edge for the Anderson Model on Random Regular Graphs

### 15 Mar 2026 v1

@@ Abstract.
We determine the phase diagram of the Anderson tight-binding model on regular random graphs with Gaussian disorder distribution and sufficiently large degree. In particular, we prove that if the degree is fixed and the number of vertices goes to infinity, the spectrum asymptotically consists of a finite delocalized interval surrounded by two unbounded localized components. Our argument uses a recent description of the spectrum of the tight-binding model on the Bethe lattice (Aggarwal–Lopatto, 2025). By viewing the Bethe lattice as the local limit of a random regular graph, and proving several preliminary estimates, we transfer this characterization of the spectrum of the limiting model to its finite approximants.
@@ 1.1. Background
A standard setting for studying wave transport in disordered media is the tight-binding model, which served as the basis for Anderson’s original investigations [12]. Let $\mathcal{G}$ be a graph (finite or infinite), let $A$ denote the adjacency operator of $\mathcal{G}$, and let $V$ denote a diagonal operator of independent, identically distributed (i.i.d) random variables. A particle whose time evolution is governed by $A$ moves freely by nearest-neighbor hopping, displaying diffusive behavior. In contrast, the time evolution induced by $V$ traps a particle at its initial site. The tight-binding Hamiltonian is defined as $H_{t}=-tA+V$; it represents a contest between the opposite influences of $A$ and $V$ mediated by $t$, the inverse disorder strength. When $t$ is sufficiently small, $V$ dominates, and particle remains frozen in position space. However, when $t$ is sufficiently large, it is expected that the particle may diffuse, at least for certain energies and graph geometries, leading to a region (or regions) of delocalization in the spectrum separated from regions of localization by sharp transitions called mobility edges [9].
@@
An important special case is when $\mathcal{G}$ is the infinite $d$-regular tree, also known as the Bethe lattice. While simpler than more physically realistic graphs, such as $\mathbb{Z}^{3}$, its lack of cycles renders it more amenable to analysis [2, 1, 29, 28]. Because of this, it has served as a distinguished testbed for theoretical investigations. For instance, the self-consistent theory of Anderson localization was developed using equations which are approximate on a true lattice, but become exact on the Bethe lattice [2, 1]. Much is now known rigorously about this model, including delocalization at small energies [27, 26, 7, 22], localization at large energies [10, 6], and a nearly complete description of the phase diagram at low disorder (omitting only small intervals around the expected locations of the mobility edges) [20]. Further, [8] established a criterion for absolutely continuous spectrum based on the fractional moments of resolvent entries. This was used in [13], along with the Simon–Wolff criterion for localization [32, 6], to provide quantitative bounds on the endpoints of the regions of localization and delocalization, mapping a large part of the spectrum. As the degree of the Bethe lattice tends to infinity, these bounds become increasingly better and characterize the entire the spectrum in the limit (but not for any finite degree). Recently, building on the work of [13, 32, 6, 8, 4], mobility edges separating regions of delocalization and localization were shown to exist in [5] when the degree $d$ is sufficiently large and the potential has unbounded support.
@@
A close cousin of the Bethe lattice is the random regular graph, which is generally understood to be its correct finite-size analog. (Simply truncating the Bethe lattice to a finite tree leads to a graph where most vertices are leaves, which produce undesirable boundary effects.) Since random regular graphs converge locally to the Bethe lattice, their spectral and eigenfunction observables measure how Bethe-lattice physics is modified by long loops and finite-size effects, providing a controlled high-dimensional reference for comparison with finite-dimensional Anderson localization [16, 17, 33]. Further, there has been a resurgence of interest in the tight-binding model on the random regular graph due to its connection with many-body localization, a phenomenon that has recently attracted significant interested in the physics community [30]. Specifically, in suitable many-body problems, certain hierarchical models in Fock space can be approximately mapped to tight-binding models on locally tree-like graphs, for which a random regular graph is a useful proxy [34].
@@
There have been only a few rigorous results for the tight-binding model on random regular graphs. In [23], it was shown that for compactly supported and bounded disorder densities, the integrated density of states converges to the integrated density of states of the Bethe lattice, as $d$ is fixed and the number of vertices $n$ tends to infinity, with a rate of convergence of order $1/\ln n$. This work also showed for that if the eigenvectors in a finite energy window in the absolutely continuous spectrum for the Bethe lattice are averaged, weighted by their squared amplitudes at a uniformly random vertex $x_{0}$, the resulting weighted mass does not concentrate in small balls around $x_{0}$; this is a form of eigenvector delocalization. Further, [24] showed a criterion for the presence of Poisson spectral statistics based on the fractional moment decay of off-diagonal resolvent entries and applied it to the edges of the spectrum (for compactly supported disorder distributions). Finally, it was shown in [11] that in the weak-disorder regime, for energy windows in the absolutely continuous spectrum for the Bethe lattice, most eigenfunctions are asymptotically equidistributed on their coordinates (again for compactly supported disorder distributions).
@@ 1.2. Main Result
Fix an integer $d\geq 3$, and let $\mathbb{N}$ denote the positive integers. For each $N\in\mathbb{N}$, let $G_{N}$ be a (uniformly) random $d$-regular graph on the vertex set $N]=\{1,\dots,N\}$, and let $A_{N}$ denote its adjacency matrix. Let $(g_{i})_{i=1}^{\infty}$ be i.i.d. Gaussian random variables with mean zero and variance one. Define the diagonal matrix
@@
where $\delta_{k\ell}$ is the Kronecker delta function. The eigenvectors $u_{k}$ are determined only up to a sign, which we may select arbitrarily. We let $\Lambda$ denote the multiset of eigenvalues of $W_{N}$ (which may have duplicated elements due to repeated eigenvalues), and given an interval $I\subset\mathbb{R}$, we let $\Lambda_{I}$ be the multiset of eigenvalues contained in $\mathbb{R}$. While these objects depend on $d$, $t$, and $N$, we suppress this dependence in the notation.
@@
To understand the meaning of these observables, we consider two extreme cases. First, suppose that all eigenvectors with eigenvalues in $I$ are as delocalized as possible, so that $u_{i}(j)=N^{-1/2}$ for all $j\in[N]$ and $i$ such that $\lambda_{i}\in I$. Then $P_{I}(j)=1$ for all $j$, implying $Q_{I}=1$. In particular, $Q_{I}$ stays bounded as $N$ grows. Next, consider the case where $u_{i}(j)=\delta_{ij}$ for all $i,j\in[N]$, in which all eigenvectors are maximally localized. Then $P_{I}(j)=0$ if $\lambda_{j}\notin I$ and $N|\lambda_{I}|^{-1}$ if $\lambda_{j}\in I$, implying $Q_{I}=N|\Lambda_{I}|^{-1}$. If $|\Lambda_{I}|\leq C|I|N$ for all $N$, where $C>0$ is a constant that does not depend on $I$ or $N$, and $|I|$ denotes the length of $I$, then $Q_{I}$ becomes unbounded as $|I|$ tends to zero.
@@
$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\leq C\Big)=1.$
@@
We now state our main theorem. Let $\rho$ denote the density function of a mean zero, variance one Gaussian random variable. The scaling of the strength disorder as $(d\ln d)^{-1}$ is conventional; it ensures the mobility edge locations tend to a constant-order limit as $d$ grows large.
@@
$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$
and let $t=g(d\ln d)^{-1}$. Then there exists $\mathfrak{E}>0$ such that the set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$. Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$, localization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|>\mathfrak{M}\}$, and delocalization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|<\mathfrak{M}\}$.
@@ 1.3. Proof Ideas
As mentioned previously, the spectral decomposition of the tight-binding model of the Bethe lattice was characterized in [5] (for the disorder regime considered in our main result). Since the Bethe lattice is the local limit of random regular graphs, it natural to attempt to use the Bethe lattice result to prove Theorem 1.3, and this is indeed our strategy. However, this route requires suitable estimates for the tight-binding model on the regular random graph, which we enumerate in Section 3. Specifically, we require concentration for fractional moments of the off-diagonal resolvent entries (Lemma 3.3), upper and lower bounds on the eigenvalue density (Lemma 3.4), convergence of the diagonal resolvent entries of the random regular graph to the diagonal resolvent entry at the root of the Bethe lattice (Lemma 3.5), and, in the delocalized regime, a bound on the second moment of the imaginary of part of the resolvent at the root of the Bethe lattice that is uniform in the imaginary part of the spectral parameter (Lemma 3.6). Of these, the last is perhaps the most technically challenging. We prove it by first establishing a Cauchy-type tail bound on the imaginary part of the relevant resolvent entry, then bootstrapping that bound to an improved polynomial tail decay bound.
Generalizations of our result are possible in at least two directions. First, we have chosen to concentrate on the case of Gaussian disorder to streamline the presentation. The spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] for a wide class of unbounded potentials; in the general case, it is slightly more complicated to describe, with interleaved regions of localization and delocalized determined by the disorder density. The techniques we present suffice to establish the analogous result for random regular graphs with corresponding potentials under the additional assumption that density decays sufficiently rapidly (for example, exponentially), so that the concentration argument for Lemma 3.3 below, based on Gaussian concentration for Lipschitz functions, can be replaced by one based on truncation and Talagrand’s concentration inequality for bounded distributions.
Second, the requirement that $d$ be very large in Theorem 1.3 comes primarily from the analogous Bethe lattice result in [5], not our arguments that show its conclusions transfer to random regular graphs. The only input we require from the limiting model is an understanding of how the imaginary part of the resolvent at the root behaves in the upper half plane as the imaginary part of its spectral parameter goes to zero (see Lemma 2.3), along with a mild lower bound on $d$ (which could be improved). In particular, it seems likely that our methods suffice to transfer the nearly complete description of the phase diagram for the Bethe lattice at low disorder recently established in [20] to random regular graphs.
@@ 2.1. Notations and Conventions
We define the adjacency operator $A=A^{(d)}$ for $\mathbb{T}$ by setting $\langle\delta_{v},A\delta_{w}\rangle=1$ if there is an edge between $v$ and $w$ in $\mathbb{T}$ and setting $\langle\delta_{v},A\delta_{w}\rangle=0$ otherwise. It extends by linearity to a bounded operator $\ell^{2}(\mathbb{V})$. Further, we let $(g_{v})_{v\in\mathbb{V}}$ be collection of i.i.d. Gaussian random variables with mean zero and variance one. We define an unbounded operator $V=V^{(d)}$ on $\ell^{2}(\mathbb{V})$ by setting $\langle\delta_{v},V\delta_{w}\rangle=g_{v}$ if $v=w$ and $\langle\delta_{v},V\delta_{w}\rangle=0$ otherwise, then extending by linearity. Then, for every $t>0$, we define the operator
@@
It is well-known that $H_{d,t}$ is unbounded and self-adjoint on the domain of $V^{(d)}$; see, e.g., [25].
@@ 2.2. Mobility Edge on the Bethe Lattice
$\frac{1}{4\|\rho\|_{\infty}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$ (2.1)
and let $t=g(d\ln d)^{-1}$. Then there exists $\mathfrak{E}>0$ such that the solution set $\{E:\rho(E)=1/(4g)\}$ equals $\{-\mathfrak{E},\mathfrak{E}\}$. Further, there exists $\mathfrak{M}(d)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathcal{L}^{-1}$ and the following claims hold.
@@
    $\lim_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R(E+\mathrm{i}\eta)>c)>c.$
@@
In [5], the conclusion of the theorem is stated in a slightly different way, involving absolutely continuous and pure point spectrum. However, it is readily verified by inspecting the proof that the claims in Theorem 2.1 are proved as intermediate steps (see, e.g., [5, Lemma 2.7].) We also note that the proof in [5] is for a tree whose root degree is one less than each other vertex. However, as mentioned in that reference, is it straightforward to see that the same argument gives the conclusion for $\mathbb{T}_{d}$.

### 19 Mar 2026 v2

@@ Abstract.
We determine the phase diagram of the Anderson tight-binding model on random regular graphs with Gaussian disorder and sufficiently large degree. In particular, we prove that if the degree is fixed and the number of vertices goes to infinity, the spectrum asymptotically consists of a finite delocalized interval surrounded by two unbounded localized components. Our argument uses a recent description of the spectrum of the tight-binding model on the Bethe lattice (Aggarwal–Lopatto, 2025). By viewing the Bethe lattice as the local limit of a random regular graph, and establishing suitable concentration, eigenvalue-counting, and resolvent estimates, we transfer this characterization of the spectrum of the limiting model to the finite-volume setting.
@@ 1.1. Background
A standard setting for studying wave transport in disordered media is the tight-binding model, which served as the basis for Anderson’s original investigations [12]. Let $\mathcal{G}$ be a graph (finite or infinite), let $A$ denote the adjacency operator of $\mathcal{G}$, and let $V$ denote a diagonal operator of independent, identically distributed (i.i.d.) random variables. A particle whose time evolution is governed by $A$ moves freely by nearest-neighbor hopping. In contrast, the time evolution induced by $V$ traps a particle at its initial site. The tight-binding Hamiltonian is defined as $H_{t}=-tA+V$; it represents a contest between the opposite influences of $A$ and $V$ mediated by $t$, the inverse disorder strength. When $t$ is sufficiently small, $V$ dominates, and the particle remains localized near its initial position. However, when $t$ is sufficiently large, one expects that extended states may emerge, at least for certain energies and graph geometries, leading to a region (or regions) of delocalization in the spectrum separated from regions of localization by sharp transitions called mobility edges [9].
@@
An important special case is when $\mathcal{G}$ is the infinite $d$-regular tree, also known as the Bethe lattice. While simpler than more physically realistic graphs, such as $\mathbb{Z}^{3}$, its lack of cycles renders it more amenable to analysis [2, 1, 30, 29]. Because of this, it has served as a distinguished testbed for theoretical investigations. For instance, the self-consistent theory of Anderson localization was developed using equations which are approximate on a true lattice, but become exact on the Bethe lattice [2, 1]. Much is now known rigorously about this model, including delocalization at small energies [26, 28, 27, 7, 22], localization at large energies [10, 6], and a nearly complete description of the phase diagram at low disorder (omitting only small intervals around the expected locations of the mobility edges) [21]. Further, [8] established a criterion for absolutely continuous spectrum based on the fractional moments of resolvent entries. This was used in [14], along with the Simon–Wolff criterion for localization [33, 6], to provide quantitative bounds on the endpoints of the regions of localization and delocalization, mapping a large part of the spectrum. As the degree of the Bethe lattice tends to infinity, these bounds become increasingly better and characterize the entire spectrum in the limit (but not for any finite degree). Recently, building on the work of [14, 33, 6, 8, 4], mobility edges separating regions of delocalization and localization were shown to exist in [5] when the degree $d$ is sufficiently large and the potential has unbounded support.
@@
A close cousin of the Bethe lattice is the random regular graph, which is generally understood to be its correct finite-size analog. (Simply truncating the Bethe lattice to a finite tree leads to a graph where most vertices are leaves, which produce undesirable boundary effects.) Since random regular graphs converge locally to the Bethe lattice, their spectral and eigenfunction observables measure how Bethe-lattice physics is modified by long loops and finite-size effects, providing a controlled high-dimensional reference for comparison with finite-dimensional Anderson localization [17, 18, 34]. Further, there has been a resurgence of interest in the tight-binding model on the random regular graph due to its connection with many-body localization, a phenomenon that has recently attracted significant interest in the physics community [31]. Specifically, in suitable many-body problems, certain hierarchical models in Fock space can be approximately mapped to tight-binding models on locally tree-like graphs, for which a random regular graph is a useful proxy [35].
@@
There have been only a few rigorous results for the tight-binding model on random regular graphs. In [23], it was shown that for compactly supported and bounded disorder densities, the integrated density of states converges to the integrated density of states of the Bethe lattice, as $d$ is fixed and the number of vertices $N$ tends to infinity, with a rate of convergence of order $1/\ln N$. This work also showed that if the eigenvectors in a finite energy window in the absolutely continuous spectrum for the Bethe lattice are averaged, weighted by their squared amplitudes at a uniformly random vertex $x_{0}$, the resulting weighted mass does not concentrate in small balls around $x_{0}$; this is a form of eigenvector delocalization. Further, [24] showed a criterion for the presence of Poisson spectral statistics based on the fractional moment decay of off-diagonal resolvent entries and applied it to the edges of the spectrum (for compactly supported disorder distributions). Finally, it was shown in [11] that in the weak-disorder regime, for energy windows in the absolutely continuous spectrum for the Bethe lattice, most eigenfunctions are asymptotically equidistributed on their coordinates (again for compactly supported disorder distributions).
@@ 1.2. Main Result
Fix an integer $d\geq 3$, and let $\mathbb{N}$ denote the positive integers. When discussing $d$-regular random graphs with $N$ vertices, we always restrict to the subset of integers $N$ such that $N\geq d+1$ and $Nd$ is even (so that the configuration model construction of $d$-regular graphs is valid). Statements quantifying over $N$ implicitly quantify over only this subset, and all limits $N\rightarrow\infty$ are implicitly along the subsequence of these values.
Let $\mathcal{G}_{N}$ be a (uniformly) random $d$-regular graph on the vertex set $N]=\{1,\dots,N\}$, and let $A_{N}$ denote its adjacency matrix. Let $(g_{i})_{i=1}^{\infty}$ be i.i.d. Gaussian random variables with mean zero and variance one (independent of $\mathcal{G}_{N}$). Define the diagonal matrix
@@
where $\delta_{k\ell}$ is the Kronecker delta function. The eigenvectors $u_{k}$ are generically determined only up to a sign, which we may select arbitrarily. We let $\Lambda$ denote the multiset of eigenvalues of $W_{N}$ (which may have duplicated elements due to repeated eigenvalues), and given an interval $I\subset\mathbb{R}$, we let $\Lambda_{I}$ be the multiset of eigenvalues contained in $I$. While these objects depend on $d$, $t$, and $N$, we suppress this dependence in the notation.
@@
When $|\Lambda_{I}|=0$, we set $P_{I}(j)=0$.
To understand the meaning of these observables, we consider two extreme cases. First, suppose that all eigenvectors with eigenvalues in $I$ are as delocalized as possible, so that $u_{i}(j)^{2}=N^{-1}$ for all $j\in[N]$ and $i$ such that $\lambda_{i}\in I$. Then $P_{I}(j)=1$ for all $j$, implying $Q_{I}=1$. In particular, $Q_{I}$ stays bounded as $N$ grows. Next, consider the case where $u_{i}(j)=\delta_{ij}$ for all $i,j\in[N]$, in which all eigenvectors are maximally localized. Then $P_{I}(j)=0$ if $\lambda_{j}\notin I$ and $N|\Lambda_{I}|^{-1}$ if $\lambda_{j}\in I$, implying $Q_{I}=N|\Lambda_{I}|^{-1}$. If $|\Lambda_{I}|\leq C|I|N$ for all $N$, where $C>0$ is a constant that does not depend on $I$ or $N$, and $|I|$ denotes the length of $I$, then $Q_{I}$ becomes unbounded as $|I|$ tends to zero.
@@
$\lim_{N\to\infty}\mathbb{P}\Big(Q_{I(\varepsilon)}\leq C\Big)=1,$
@@
We now state our main theorem. The scaling of the inverse disorder strength as $(d\ln d)^{-1}$ is conventional; it ensures the mobility edge locations tend to constant-order limits as $d$ grows large. The constants arise from the fact that the supremum of the density of a mean zero, variance one Gaussian random variable is $(2\pi)^{-1/2}$.
@@
$\sqrt{\frac{\pi}{8}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$
and let
$t=g(d\ln d)^{-1},\qquad\mathfrak{E}=\sqrt{2\ln\!\left(\frac{4g}{\sqrt{2\pi}}\right)}.$
Then there exists $\mathfrak{M}(d,g)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathfrak{L}^{-1}$, localization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|>\mathfrak{M}\}$, and delocalization holds for $W_{N}$ at all values in the set $\{E\in\mathbb{R}:|E|<\mathfrak{M}\}$.
@@ 1.3. Proof Ideas
As mentioned previously, the spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] (for the disorder regime considered in our main result). Since the Bethe lattice is the local limit of random regular graphs, it is natural to attempt to use the Bethe lattice result to prove Theorem 1.3, and this is indeed our strategy. However, this route requires suitable estimates for the tight-binding model on the regular random graph, which we enumerate in Section 3. Specifically, we require concentration for fractional moments of the diagonal resolvent entries (Lemma 3.3), upper and lower bounds on the eigenvalue density (Lemma 3.5), convergence of the diagonal resolvent entries of the random regular graph to the diagonal resolvent entry at the root of the Bethe lattice (Lemma 3.4), and, in the delocalized regime, a bound on the second moment of the imaginary part of the resolvent at the root of the Bethe lattice that is uniform in the imaginary part of the spectral parameter (Lemma 3.6). Of these, the last is perhaps the most technically challenging. We prove it by first establishing a Cauchy-type tail bound on the imaginary part of the relevant resolvent entry, then bootstrapping that bound to an improved polynomial tail decay bound.
Generalizations of our result are possible in at least two directions. First, we have chosen to concentrate on the case of Gaussian disorder to streamline the presentation. The spectral decomposition of the tight-binding model on the Bethe lattice was characterized in [5] for a wide class of unbounded potentials; in the general case, it is slightly more complicated to describe, with interleaved regions of localization and delocalization determined by the disorder density. The techniques we present suffice to establish the analogous result for random regular graphs with corresponding potentials under the additional assumption that the density decays sufficiently rapidly (for example, exponentially), so that the concentration argument for Lemma 3.3 below, based on Gaussian concentration for Lipschitz functions, can be replaced by one based on truncation and Talagrand’s concentration inequality for bounded distributions.
Second, the requirement that $d$ be very large in Theorem 1.3 comes primarily from the analogous Bethe lattice result in [5], not our arguments that show its conclusions transfer to random regular graphs. The main input we require from the limiting model is an understanding of how the imaginary part of the resolvent at the root behaves in the upper half plane as the imaginary part of its spectral parameter goes to zero (see Lemma 2.3). In particular, it seems likely that our methods suffice to transfer the nearly complete description of the phase diagram for the Bethe lattice at low disorder recently established in [21] to random regular graphs.
@@ 2.1. Notations and Conventions
We define the adjacency operator $A=A^{(d)}$ for $\mathbb{T}$ by setting $\langle\delta_{v},A\delta_{w}\rangle=1$ if there is an edge between $v$ and $w$ in $\mathbb{T}$ and setting $\langle\delta_{v},A\delta_{w}\rangle=0$ otherwise. It extends by linearity to a bounded operator $\ell^{2}(\mathbb{V})$. Further, we let $(g_{v})_{v\in\mathbb{V}}$ be a collection of i.i.d. Gaussian random variables with mean zero and variance one. We define an unbounded operator $V=V^{(d)}$ on $\ell^{2}(\mathbb{V})$ by setting $\langle\delta_{v},V\delta_{w}\rangle=g_{v}$ if $v=w$ and $\langle\delta_{v},V\delta_{w}\rangle=0$ otherwise, then extending by linearity. Then, for every $t>0$, we define the operator
@@
It is well-known that $H_{d,t}$ is unbounded and self-adjoint on the domain of $V^{(d)}$; see, e.g., [25]. For all $z\in\mathbb{H}$, we let $R(z)$ denote the resolvent of $H_{d,t}$.
@@ 2.2. Mobility Edge on the Bethe Lattice
$\sqrt{\frac{\pi}{8}}+\frac{1}{\mathfrak{L}}<g<\mathfrak{L},$ (2.1)
and let
$t=g(d\ln d)^{-1},\qquad\mathfrak{E}=\sqrt{2\ln\!\left(\frac{4g}{\sqrt{2\pi}}\right)}.$
Then there exists $\mathfrak{M}(d,g)>0$ such that $|\mathfrak{M}-\mathfrak{E}|<\mathfrak{L}^{-1}$ and the following claims hold.
@@
    $\liminf_{\eta\rightarrow 0}\mathbb{P}(\operatorname{Im}R_{00}(E+\mathrm{i}\eta)>c)>c.$
@@
In [5], the conclusion of the theorem is stated in a slightly different way, involving absolutely continuous and pure point spectrum. However, it is readily verified by inspecting the proof that the claims in Theorem 2.1 are proved as intermediate steps (see, e.g., [5, Lemma 2.7]). We also note that the proof in [5] is for a tree whose root degree is one less than each other vertex. However, as mentioned in that reference, it is straightforward to see that the same argument gives the conclusion for $\mathbb{T}_{d}$.